carmelle:

Choose the correct simplification of (9x3y2)2. 18x5y4 81x5y4 18x6y4 81x6y4

XioGonz:

\[(9x3y2)^2\]^^ Like that?

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XioGonz \[(9x3y2)^2\]^^ Like that? \(\color{#0cbb34}{\text{End of Quote}}\) yea, mb

XioGonz:

\[(9x3y2)^2\] \[(3^2 x^3 • (y^2))^2\] x^3 raised to the 2nd power = x^( 3 * 2 ) =x^6 y^2 raised to the 2nd power =y^( 2 * 2 ) =y^4 HENCE, your answer is \[3^4x^6y^4\]

carmelle:

That's not one of the answer choices XD

xxDeppressionxx:

Now i am confused even more what is the answer and what isn't the answer?

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx Now i am confused even more what is the answer and what isn't the answer? \(\color{#0cbb34}{\text{End of Quote}}\) 18x5y4 81x5y4 18x6y4 81x6y4 those are the answer choices ^^^ 3^4x^6y^4 that one isn't ^

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sealake101 i think its 81x6y4 \(\color{#0cbb34}{\text{End of Quote}}\) how tho, can u explain?

xxDeppressionxx:

Make sure to always explain your answer and how you got it

carmelle:

I'll just close this :/

sealake101:

9^2 * (x^3)^2 * (y^2)^2 9 squared is 81: 81 * (x^3)^2 * (y^2)^2 (x^3)^2 = x^(3*2) = x^6 (y^2)^2 = y^(2*2) = y^4 them having those from ur choices its 81(x^6)(y^4)

sealake101:

hope that helps :)